OpenStudy (anonymous):

No one has been able to help! Please help I'll give a medal! Simplify the rational expression x^2+2x-35/x^2+4x-21

OpenStudy (anonymous):

@jim_thompson5910 do you mind helping me please?!

jimthompson5910 (jim_thompson5910):

Are you able to factor any of the expressions?

OpenStudy (anonymous):

The answer I got was x+5/x+3 but idk if that is right

jimthompson5910 (jim_thompson5910):

what do you get when you factor the numerator?

OpenStudy (anonymous):

I got x^2+2x-35=(x+7)(x-5)

jimthompson5910 (jim_thompson5910):

how about the denominator

OpenStudy (anonymous):

(x-7)(x+3)

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

notice how -7+3 = -4 but we want +4 instead

OpenStudy (anonymous):

Oh ok! so what do I do next

jimthompson5910 (jim_thompson5910):

what's the correct factorization of the denominator?

OpenStudy (anonymous):

7-3=4 so (x+7)(x-3)?

jimthompson5910 (jim_thompson5910):

So we will get this: \[\Large \frac{x^2+2x-35}{x^2+4x-21}\] \[\Large \frac{(x+7)(x-5)}{x^2+4x-21}\] \[\Large \frac{(x+7)(x-5)}{(x+7)(x-3)}\] \[\Large \frac{\cancel{(x+7)}(x-5)}{\cancel{(x+7)}(x-3)}\] \[\Large \frac{x-5}{x-3}\]

OpenStudy (anonymous):

Oh ok! Thank you for explaining it!

jimthompson5910 (jim_thompson5910):

you're welcome